# NeSy is alive and well: A LLM-driven symbolic approach for better code comment data generation and classification

### v1

@@ 3 Methodology
This section describes our NeSy methodology combining a LLM agent and a symbolic framework to generate synthetic labeled code comment data. We chose ChatGPT 3.5 to implement our methods and experiments since it is freely accessible and usable without prior configuration. We introduce a set of rules based on semantic decomposition to prompt ChatGPT and create a neuro-symbolic workflow that teaches the LLM the proper syntax of the C programming language for controlling the generation of synthetic labeled code comment samples. The workflow is represented in Figure 1.
@@ 3.1 Semantic rules
The rules themselves are renditions of the syntax of the C programming language [26] and delimit the vocabulary as well as the constructs of the language. They start at the atomic level by defining what a valid token in the language is and move to more complex concepts like determining the construction of a valid line of code in C. Each rule is written as a statement in natural language and is kept as simple and short as possible. Figure 2 shows the 12 rules given as a prompt for ChatGPT to produce a valid line of C code.
@@
1
The smallest individual unit of a program is called a token.
2
Tokens are either keywords, identifiers or variables.
3
A keyword must belong to the list: auto, double, int, struct, break, else, long, switch, case, enum, register, typedef, char, extern, return, union, const, float, short, unsigned, continue, for, signed, void, default, goto, sizeof, volatile, do, if, static, while.
4
An identifier can only have alphanumeric characters(a-z , A-Z , 0-9) and underscore(_).
5
The first character of an identifier can only contain alphabet(a-z, A-Z) or underscore (_).
6
Identifiers are case-sensitive in the C language. For example, name and Name will be treated as two different identifiers.
7
Keywords are not allowed to be used as Identifiers.
8
No special characters, such as a semicolon, period, whitespaces, slash, or comma are permitted to be used in or as an Identifier. Example of valid identifiers: total, avg1, difference_1. Example of invalid identifiers: \$myvar, x!y.
9
A variable has a data type (which can be one of the following: char, int, float, double, void), a name and a value.
10
A variable should be declared and assigned a value. Example: int marks = 10.
11
After creation and assignment, the value of a variable can be changed.
12
A valid line of code is a collection of tokens that adhere to the above rules.
13
Comments are plain simple text in English that can be added to a line of code.
14
A comment explains various parts of the line of code, makes it more readable and more understandable.
15
A comment either begins with // if it is a single-line comment or is enclosed within /* and */ if it is a multi-line comment.
16
Comments can be either labeled Useful or Not Useful.
17
A comment is labeled Useful when it is informative and helps clarify the line of code without being redundant, otherwise, it is labeled Not Useful.

### v2

@@ 3 Methodology
This section describes our NeSy methodology combining a LLM agent and a symbolic framework to generate synthetic labeled code comment data as shown in Figure 1. We chose ChatGPT 3.5 to implement our methods and experiments since it is freely accessible and usable without prior configuration. We introduce a set of rules based on semantic decomposition to prompt ChatGPT and create a neuro-symbolic workflow that teaches the LLM the proper syntax of the C programming language for controlling the generation of synthetic labeled code comment samples. The workflow is represented in Figure 2.
Figure 1: Example of labeled code comment data
@@ 3.1 Semantic rules
The rules themselves have been designed as renditions of the syntax of the C programming language [26] and delimit the vocabulary as well as the constructs of the language. They start at the atomic level by defining what a valid token in the language is and move to more complex concepts like determining the construction of a valid line of code in C. Each rule is written as a statement in natural language and is kept as simple and short as possible. Figure 3 shows the 12 rules given as a prompt for ChatGPT to produce a valid line of C code.
@@
1 The smallest individual unit of a program is called a token.
2 Tokens are either keywords, identifiers or variables.
3 A keyword must belong to the list: auto, double, int, struct, break, else, long, switch, case, enum, register, typedef, char, extern, return, union, const, float, short, unsigned, continue, for, signed, void, default, goto, sizeof, volatile, do, if, static, while.
4 An identifier can only have alphanumeric characters(a-z , A-Z , 0-9) and underscore(_).
5 The first character of an identifier can only contain alphabet(a-z, A-Z) or underscore (_).
6 Identifiers are case-sensitive in the C language. For example, name and Name will be treated as two different identifiers.
7 Keywords are not allowed to be used as Identifiers.
8 No special characters, such as a semicolon, period, whitespaces, slash, or comma are permitted to be used in or as an Identifier. Example of valid identifiers: total, avg1, difference_1. Example of invalid identifiers: \$myvar, x!y.
9 A variable has a data type (which can be one of the following: char, int, float, double, void), a name and a value.
10 A variable should be declared and assigned a value. Example: int marks = 10.
11 After creation and assignment, the value of a variable can be changed.
12 A valid line of code is a collection of tokens that adhere to the above rules.
13 Comments are plain simple text in English that can be added to a line of code.
14 A comment explains various parts of the line of code, makes it more readable and more understandable.
15 A comment either begins with // if it is a single-line comment or is enclosed within /* and */ if it is a multi-line comment.
16 Comments can be either labeled Useful or Not Useful.
17 A comment is labeled Useful when it is informative and helps clarify the line of code without being redundant, otherwise, it is labeled Not Useful.